OpenStudy (anonymous):

An isosceles trapezoid has bases of 4 and 10. If the base angle is 45°, find the area. help me

OpenStudy (anonymous):

if u drop a perpendicular line from end of the side length 4 you get a triangle with base (10-4)/2 = 3 so height = 3 tan45 = 3 area = (1/2) * 3 * (4+10) = 21 sq units

OpenStudy (anonymous):

cherri - where are u?